### v3

# Learning New Tasks from a Few Examples with Soft-Label Prototypes

Abstract
Existing approaches to few-shot learning in NLP rely on large language models and fine-tuning of these to generalise on out-of-distribution data. In this work, we propose a simple yet powerful approach to “extreme” few-shot learning, wherein models are exposed to as little as $4$ examples per class, based on soft-label prototypes that collectively capture the distribution of different classes across the input domain space. Inspired by previous work Sucholutsky et al. (2021) on univariate or simple multivariate (synthetic) data, we propose a novel approach that is effective on large, high-dimensional and real-world datasets. We learn soft-label prototypes within a neural framework (DeepSLP) and we experimentally demonstrate that it achieves superior performance on 31/48 tested tasks and few-shot settings while closely matching the performance of strong baselines on the rest. We focus on learning previously unseen NLP tasks from very few examples ($4$, $8$, $16$) and present an in-depth analysis of the effectiveness of our approach.
1 Introduction
Humans have a remarkable ability to adapt knowledge gained in one domain and apply it in another setting, and to identify or disambiguate objects after observing only a handful of examples Lake et al. (2015). This has inspired research in few-shot learning that aims to build models that can also learn a new task using only a small number of examples per class. Early few-shot learning in NLP relied on interventions at the data level, such as dataset augmentation Clark et al. (2018) or generation of adversarial examples from few-shot datasets Miyato et al. (2016). The advent of large language models (LLMs) has now led to a range of approaches, including fine-tuning on different target tasks Sun et al. (2020); Zhou and Srikumar (2022), creating prompt-enhanced few-shot datasets for training Gao et al. (2020); Schick and Schütze (2020); Lester et al. (2021), and the application of meta-learning Finn et al. (2017); Snell et al. (2017) to optimise model parameters such that models adapt quickly to new tasks using past experience Dou et al. (2019); Holla et al. (2020); van der Heijden et al. (2021) which requires a large number of diverse and balanced few-shot meta-training tasks for effective learning.
In this paper, we propose a simple and effective approach to “extreme” few-shot learning where a model generalises on unseen data after adaptation on as little as 4 examples per class, in the absence of any auxiliary training data and without any LLM finetuning. Inspired by Sucholutsky et al. (2021), we first generate compact representations of input data (by solving for linear constraints) called soft-label prototypes that capture the distribution of different classes across the input domain space, and which are used to make predictions. While the original method of Sucholutsky et al. (2021) is designed for small sets of simple univariate or multivariate (synthetic) data, we extend their approach to work on large sets of high-dimensional and real-world NLP datasets. We then propose a novel method that learns such soft-label prototypes within a neural framework, DeepSLP. We focus on learning previously unseen NLP tasks from a few examples per task ($4$, $8$, $16$). We experimentally demonstrate that DeepSLP outperforms a range of strong few-shot baselines in the majority ($31/48$) of investigated tasks, while closely matching their performance on the remaining tasks.
2 Related work
Few-shot learning approaches in NLP include data augmentation and semi-supervised learning; e.g., augmentation with adversarial examples Miyato et al. (2016), interpolation of training data into a learnable higher dimensional embedding space Chen et al. (2020), and consistency training to make models more resistant to noise Xie et al. (2019). Recent research efforts on large-scale pre-training of language models Devlin et al. (2018); Radford et al. (2019); Brown et al. (2020) reduce the amount of data required for their subsequent fine-tuning in a given task. Instruction tuning Sanh et al. (2021) and in-context learning Liu et al. (2021) show that natural language instructions or prompts can enhance a model’s few-shot learning abilities by leveraging the language (instruction) understanding abilities of the given pretrained LLM. Creating masked language modeling tasks (at training time) with prompts, either in the vocabulary space Gao et al. (2020) or embedding space Li and Liang (2021); Lester et al. (2021), is experimentally demonstrated to give improved performance across classification tasks in few-shot settings compared to LLM fine-tuning for GPT2 Radford et al. (2019) and BART Lewis et al. (2019). Sanh et al. (2021) further report that diverse and additional multi-task prompted training leads to better generalisation, less variation on unseen data and an overall superior performance compared to larger state-of-the-art models. However, the search space over LLMs, prompt templates, and few-shot learning is so great that there is yet to be an established standard.
Previous work has also tackled few-shot learning within the meta-learning paradigm of “learning to learn” , referring to methods that are trained to adapt quickly (in a few gradient steps) and from a small number of examples using past experience Finn et al. (2017). Meta-learning has been applied to a range of tasks, including relation extraction and relation classification Obamuyide and Vlachos (2019a, b); word sense disambiguation Holla et al. (2020); sentiment classification, entity-typing and natural language inference Bansal et al. (2020); Nooralahzadeh et al. (2020); part-of-speech tagging Wang et al. (2020); and dependency parsing Langedijk et al. (2021).
Despite the many successful applications, however, meta-learning comes with complexities with respect to data and computation Antoniou et al. (2019). Such methods typically require a large set of diverse and balanced meta-learning tasks for effective learning Holla et al. (2020), which presents an experimental obstacle to their widespread use, while training times can span several days as their optimisation is computationally expensive Bansal et al. (2020).
Our work is more similar to the few-shot learning approach of Bansal et al. (2019), particularly in terms of tasks and datasets used. However, their approach utilises meta-learning together with auxiliary training data (GLUE; Wang et al. (2018)) and it is therefore not directly comparable to ours that focuses on a more realistic (yet more challenging) few-shot learning setting that does not rely on any additional labelled data. To facilitate a comparison, however, we do present an analysis of model performance when such data is available (Section 9.2).
3 Soft-Label Prototypes (SLP)
Sucholutsky and Schonlau (2021) define a soft label as a vector $Y$ representing a data point’s simultaneous membership to several classes. It is used in cases where we can denote a point’s partial association to different classes. Using this definition, they further define a soft-label prototype as $(\vec{X},Y)$, where $\vec{X}$ is a point in input space (e.g., an input feature vector) and $Y$ is the corresponding soft label. The underlying idea is that a small set of soft-label prototypes can be used to accurately represent a training set.
3.1 Generating soft-label prototypes
Soft-label prototypes assign soft labels to every point in the input domain; therefore, a soft-label prototype at point $\vec{X}$ represents the class distribution (determined from the training data) at $\vec{X}$. Based on Algorithm 4 in Sucholutsky et al. (2021), the process of generating soft-label prototypes from training data is a two step process — (1) finding lines that connect the class centroids in the data, where each line connects some of the centroids, and every centroid belongs to one line; and (2) deriving soft-label prototypes capturing the class distribution at the ends of each line.
3.1.1 Finding lines connecting all centroids
The underlying idea here is that we seek to find classes that lie on the same manifold. First, we compute the centroid of each class in the input data. Then, we find and fit class centroids on the minimum number of lines. For this, we use the recursive regression method proposed by Sucholutsky et al. (2021). This method clusters centroids hierarchically to group similar (interval) centroids together, and fits a regression line on the centroids. The similarity of centroids within a single cluster is judged by how well all the centroids fit on a regression line. If the Euclidean distance of a particular centroid is beyond a pre-defined tolerance threshold $\epsilon$ from a line, it is removed from that cluster and assigned to another cluster. We use this method for all our experiments, as we experimentally find (on our development data) that it performs well on high-dimensional data spread across many classes such as the ones we test here. In Figure 0(a), we present an example set of lines connecting all centroids. For further details on recursive regression, we refer the reader to Sucholutsky et al. (2021).
Figure 1: Generating and classifying data with soft-label prototypes.
3.1.2 Deriving soft-label prototypes by optimising for linear constraints
Once we find the lines, we use the endpoints of each line as the location of soft-label prototypes. Therefore, for $l$ lines fitted on $n$ centroids we have $2l$ prototypes. Sucholutsky et al. (2021) mathematically define finding the class distribution at each end point / soft-label prototype as a constraint optimisation problem. The optimisation problem consists conceptually of two main sets of constraints: (i) the target class at each centroid has the maximum influence amongst all classes at certain points along the line (endpoint of the line and midpoints between classes); and (ii) the difference between the influence of the target class and the sum of the influences of all other classes along the line is maximised. The output of these optimisation constraints are two soft-label prototypes which are “set” at the ends of each line (i.e., $\vec{X}$) as shown in Figure 0(b). For further details, we refer the reader to Sucholutsky et al. (2021).
3.2 Classification with soft-label prototypes
Given $M$ soft-label prototypes representing the input distribution of $N$ classes, we define $S=(\vec{X_{1}},Y_{1}),...,(\vec{X_{M}},Y_{M})$ to be our set of prototypes, where $\vec{X_{i}}$ is the location of the $i^{th}$ prototype in the input feature space and $Y_{i}$ is a matrix of dimension $[N\times 1]$ denoting the soft labels. Given a test datapoint $\vec{x}$, we find the nearest line to the test point and calculate the Euclidean distances ${D}={(\vec{X_{i}},\vec{x})}_{i=1,2...M}$ from each of the two prototypes of that line to $\vec{x}$. We then sort $S$ in ascending order of distances using ${D}$ and weigh the probability distribution of the $i^{th}$ nearest prototype (with a total of two prototypes) inversely by its distance to the point located at $\vec{x}$ to get $Y^{*}$:
$Y^{*}=\sum_{i=1}^{2}\frac{Y_{i}}{d(\vec{X_{i}},\vec{x})}$ (1)
[…]
3.3 Classification with SLP: A toy example
Figure 2 presents an example classification with SLP. Given the class centroids for blue, green and yellow are located at $(0,0)$, $(1.5,0)$ and $(3,0)$ respectively, two soft-label prototypes are defined by a line connecting yellow and blue, and are thus located at $(3,0)$ and $(0,0)$ respectively. The soft labels in Figure 1(a) contain the per-class probability distribution derived by the SLP method; for example, $p(x=blue)=0.6$ and $p(x=green)=0.4$ for the left prototype, and $p(x=green)=0.4$ and $p(x=yellow)=0.6$ for the right prototype. When a new test instance $x$ located at $(1.5,0.8)$ is presented, we make predictions as follows: we find the nearest line to $x$ and consider its distance from the two prototypes at the ends of the line and multiply the class distribution of each prototype by the inverse distance as per Eq. 1.
Figure 2: Classification example with SLP (figure from Sucholutsky and Schonlau (2021)).
Since $x$ is equidistant from both prototypes, the distance between the $x$ and each prototype is $1.5$. Therefore, the values for blue and yellow (given both soft-label prototypes) become $soft\_label(x=blue)=\frac{0.6}{1.5}+\frac{0}{1.5}=0.4$ and $soft\_label(x=yellow)=\frac{0}{1.5}+\frac{0.6}{1.5}=0.4$ respectively. In contrast, for green, which is directly informed by both prototypes (i.e., no zero values in the numerator), the probability distribution becomes $soft\_label(x=green)=\frac{0.4}{1.5}+\frac{0.4}{1.5}=0.53$. Therefore $x$ is classified as green. This decision boundary can be seen in Figure 1(b).
4 Methodology
We encode training and test instances (Section 5) using base cased BERT Devlin et al. (2018) from the HuggingFace Transformers [^1] https://huggingface.co/docs/transformers/index library Wolf et al. (2019). We use the training data to generate soft-label prototypes, and the test data to make predictions as per Eq. 1. We develop and experiment with two variants of the SLP approach, which we describe below. We release our code and models to facilitate further research in the field.
4.1 SLP for NLP data
We use the SLP approach described in Section 3. In order to make SLP powerful enough for large, high-dimensional real-world NLP data (vs univariate, simple multivariate and/or synthetic data), we require an optimiser that scales on such complex data. To this end, we use the MOSEK solver for linear programming MOSEK ApS (2019) in the CVXPY library Diamond and Boyd (2016) to perform the required computations to generate the soft-label prototype class distributions.
4.2 DeepSLP: Learning soft labels via gradient descent
We further develop a novel gradient-based approach to generate soft labels as a function of an input $x$ by minimising training loss on a few-shot dataset. After generating lines connecting all class centroids (Section 3.1.1), we set two soft-label prototypes at the ends of the lines. Each soft-label prototype $p_{i}$ is denoted by $g_{\theta_{i}}(f(x))$ where $g$ is a neural network parameterised by $\theta_{i}$ and $f(x)$ is a point in the input space. The neural network consists of a fixed encoder (BERT) given by $f(x)$, and a trainable linear layer which returns the soft-label probability distribution at any point $x$ given by $g_{\theta_{i}}(x)$. Figure 3 presents a visual representation of our approach. Compared to the “standard” SLP variant (Section 4.1) where we find soft-label probability distributions via linear constraint optimisation, we now parameterise our soft-label probability distribution with a neural network.
Figure 3: Learning soft-label prototypes (soft labels in red) using a trainable linear layer (yellow): example for a 3-class prototype.
The trainable linear layer is trained in a few-shot manner (see data setting in Section 5). Crucially, the encoder parameters are frozen as we need our input data points to have an unchanged location in the input space – changing their position might result in class centroids that were previously lying on a straight line to no longer lie on the line. At prediction time, we utilise both soft-label prototypes along the line using Equation 1. A higher distance between a data point $x$ and a prototype leads to a correspondingly smaller effect of the prototype on the final classification; therefore, we want to penalise the prototype that is closer to $x$ more if there is an incorrect classification. Each prototype is therefore assigned a fraction of the total loss that is proportional to the other prototype’s Euclidean distance from $x$. This way, the closer prototype’s weights are corrected more in case of a misclassification. Our approach (DeepSLP) is outlined in Algorithm 1, while an example optimisation is presented in Figure 4.
Figure 4: Training soft-label prototypes in DeepSLP. Class centroids are represented with large circles that lie on a line (Red, Green, Blue), while training set examples are represented with smaller circles of the same colour. Dotted lines represent the backpropagation error, of which the bolded ones represent a larger error per soft-label prototype (in case of misclassification). Predictions for $x$ are based on the prototypes at each end of the line.
We use cross entropy loss which gives a measure of the difference between the true and predicted labels. We initialise the weights of $f_{\theta}(x)$ using a uniform Xavier initialisation Glorot and Bengio (2010) and use warmup steps to adjust the learning rate. Epochs vary based on the number of classes in the classifier head (between 15 and 25; see datasets used in Section 5): preliminary experiments on the development data show that more epochs are needed when a higher number of classes lie along a line.
1 $\lambda\leftarrow$ set of lines connecting all centroids
2 $f_{\theta_{l}}$ is the network parameterised by $\theta_{l}$ for the left-end prototype on a line
3 $f_{\theta_{r}}$ is the network parameterised by $\theta_{r}$ for the right-end prototype on a line
4 $\mathcal{J}\leftarrow$ loss function
5 $D\leftarrow$ training data
[…]
7 for $i\in\lambda$ do
[…]
9             $p_{il}\leftarrow$ location of left prototype
10             $p_{ir}\leftarrow$ location of right prototype
11             for $x\in minibatch(D)$ do
12                   $d_{1}\leftarrow$ $||p_{il}-x||$
13                   $d_{2}\leftarrow$ $||p_{ir}-x||$
14                   $pred.append\left(\frac{f_{\theta_{il}}(x)}{d_{1}}+\frac{f_{\theta_{ir}}(x)}{d_{2}}\right)$
15             end for
16            $d\_loss\leftarrow\mathcal{J}(pred,D)$
17             $loss_{1}\leftarrow\frac{d_{2}}{d_{1}+d_{2}}*d\_loss$
18             $loss_{2}\leftarrow\frac{d_{1}}{d_{1}+d_{2}}*d\_loss$
19             $\theta_{il}\leftarrow\theta_{il}-\eta\nabla_{\theta_{il}}loss_{1}$
20             $\theta_{ir}\leftarrow\theta_{ir}-\eta\nabla_{\theta_{ir}}loss_{2}$
21       end for
22
23 end for
Algorithm 1 DeepSLP
5 Tasks and datasets
We use the same tasks and datasets as in Bansal et al. (2019). These cover a variety of text classification tasks with a large number of classes, and include: (a) Entity typing – the CoNLL-2003 Sang and De Meulder (2003) and MIT-Restaurant Liu et al. (2013) datasets; (b) Review rating classification – review ratings from Amazon Reviews Blitzer et al. (2007) with a three-way classification; (c) Text classification – scraped social media data from crowdflower [^2] https://www.figure-eight.com/data-for-everyone/ comprising sentiment and emotion classification in a range of domains, as well as political bias detection; and (d) Natural language inference in the scientific domain – the SciTail dataset Khot et al. (2018). We randomly sample $1000$ instances from the emotion task in (c) and use them as our held-out validation set for hyperparameter tuning.
We use the same data splits as Bansal et al. (2019), which are publicly available. [^3] https://github.com/iesl/leopard/ At evaluation time, Bansal et al. (2019) sample an N-way k-shot training (support set) subset $\forall\;k\in\{4,8,16\}$ for few-shot fine-tuning and then evaluate model performance on the test set of each respective task. As performance is dependent on the data used for fine-tuning, they sample $10$ different subsets per task, and report the averaged performance across subsets on the test set. We utilise the same subsets as Bansal et al. (2019) for fine-tuning our models too. Unlike Bansal et al. (2019), however, we do not utilise auxiliary training data (GLUE; Wang et al. (2018)), and only train our models on the few-shot training subsets provided. This makes our setting more challenging but one which is testing a more realistic few-shot setting that does not rely on additional data and labels. While we are not directly comparable to Bansal et al. (2019), to facilitate such a comparison, we present an analysis where we also utilise GLUE data (Section 9.2).
Similar to Bansal et al. (2019), we use mean accuracy and standard deviation to evaluate performance. For entity typing tasks, we obtain token-level encodings for classification.
6 Baselines
Our aim is to determine how well our soft-label prototype models perform compared to other classifiers given the same few-shot training data. Again, our focus is on performing an extreme few-shot learning scenario where no auxiliary data is available, and so we choose such baseline comparisons accordingly.
1-Nearest Neighbour Classifier
Our first baseline is a 1-Nearest Neighbour classifier for each task. This is motivated by the fact that test time predictions using soft-label prototypes depend on the soft-label functions of the closest class centroids to a test point $x$ (Section 3.2). We first encode all data points using BERT, and then assign test instances the class of their closest training example in this embedding space based on Euclidean distance.
Fine-tuned BERT
We use the BERT baseline reported in Bansal et al. (2019), which is fine-tuned (all layers) on the training set of each task and is trained using the same experimental/few-shot setting as ours. We refer to this model as BERTfine-tuned.
Prototypical Networks
We use Prototypical networks (ProtoNets) Snell et al. (2017) as another baseline to evaluate generalisation within a few-shot meta-learning setup. ProtoNet is a metric-based meta-learning method that uses an embedding function $f_{\theta}$ to encode training samples and compute a high-dimensional vector $\mu_{c}$ that is the arithmetic mean of the training data points of class $c$. It then uses a distance function $d$ to compute the similarity between a test instance $x$ and the mean vector of each class, and get the class probability distribution as:
$p(y=c|x)=softmax(-d(f_{\theta}(x),\mu_{c}))$
$=\frac{exp(-d(f_{\theta}(x),\mu_{c}))}{\sum_{c^{\prime}\in C}exp(-d(f_{\theta}(x),\mu_{c}^{\prime}))}$
The model is optimised using negative log-likelihood. We use BERT as the encoder and fine-tune it (all layers) using our few-shot training sets and a learning rate of 3e-5.
7 Experimental setup
Generating lines
The hyperparameters used to generate lines are: (a) $\epsilon$, which is a control factor used to denote the maximum tolerance between a centroid and the line assigned to it using Euclidean distance—we use a tolerance value of $11e-$ 1; and (b) $l$, which denotes the maximum number of lines used to connect all centroids. We experiment with a range of values ($l\in\{0.25n,0.5n,0.75n,n-1\}$, where $n$ is the number of centroids), but find $l=\lceil n/2\rceil$ to give the best accuracy on the validation data with the minimum number of lines required. [^4] The right choice of hyperparameters is key as the optimisation process fails when it is not possible to connect $n$ centroids with $l$ lines.
DeepSLP
For DeepSLP, we find that more epochs are needed to train models with a higher number of soft labels (i.e., a higher number of classes in the output of the classifier head). In other words, 3 classes fitted on a line need more epochs compared to 2 classes fitted on a line. We use AdamW Loshchilov and Hutter (2017) as our optimiser and perform hyperparameter tuning on the validation set. We find we only need a few epochs to generalise well – these range from 5 to 10 depending on the training task. We evaluate our approach and all baselines in the following manner: we fix a random seed, train our models, and evaluate performance on the test tasks. We repeat this process across three different seeds, and report the mean and standard deviation at test time.
8 Results and discussion
Category (Classes) Shot 1NN ProtoNet BERTfine-tuned* SLPBERT DeepSLPBERT
Rating DVD (3) 4 38.289 ± 6.846 37.665 ± 7.184 32.22 ± 08.72 32.298 ± 16.263 39.566 ± 5.086
8 37.598 ± 5.364 37.008 ± 5.118 36.35 ± 12.50 32.644 ± 16.016 38.788 ± 4.449
16 37.461 ± 4.506 39.123 ± 6.004 42.79 ± 10.18 35.587 ± 17.445 40.53 ± 4.375
Rating Electronics (3) 4 37.001 ± 5.164 33.696 ± 5.55 39.27 ± 10.15 35.188 ± 16.211 39.977 ± 5.959
8 37.646 ± 4.223 37.297 ± 5.938 28.74 ± 08.22 29.624 ± 12.876 41.926 ± 3.985
16 38.633 ± 3.103 43.825 ± 5.946 45.48 ± 06.13 29.836 ± 12.753 44.917 ± 3.164
Political Bias (2) 4 52.245 ± 4.348 51.15 ± 2.454 54.57 ± 5.02 53.447 ± 3.281 53.251 ± 4.042
8 54.568 ± 3.015 56.568 ± 4.228 56.15 ± 3.75 55.824 ± 3.725 58.209 ± 5.198
16 55.884 ± 2.436 59.183 ± 4.706 60.96 ± 4.25 58.277 ± 4.128 61.479 ± 2.974
Rating Kitchen (3) 4 36.769 ± 7.738 35.914 ± 6.678 34.76 ± 11.20 28.253 ± 15.907 39.624 ± 6.787
8 37.435 ± 6.347 38.46 ± 11.124 34.49 ± 08.72 24.397 ± 11.961 41.081 ± 6.777
16 38.047 ± 3.226 46.546 ± 8.394 47.94 ± 08.28 31.926 ± 18.29 45.801 ± 4.562
Political Audience (2) 4 51.827 ± 1.754 50.976 ± 1.84 51.02 ± 1.72 51.305 ± 2.68 51.741 ± 2.827
8 53.113 ± 2.314 52.022 ± 3.964 52.80 ± 2.72 53.104 ± 3.669 54.506 ± 3.274
16 53.287 ± 1.874 54.024 ± 3.071 58.45 ± 4.98 53.888 ± 3.305 56.956 ± 3.045
Sentiment Kitchen (2) 4 60.17 ± 3.197 55.24 ± 3.427 56.93 ± 7.10 61.96 ± 4.594 60.76 ± 4.426
8 59.82 ± 2.703 62.28 ± 5.103 57.13 ± 6.60 64.83 ± 3.983 65.733 ± 3.198
16 61.85 ± 2.65 66.9 ± 5.441 68.88 ± 3.39 68.21 ± 3.298 69.18 ± 2.589
Disaster (2) 4 53.629 ± 9.391 51.474 ± 8.848 55.73 ± 10.29 52.77 ± 10.803 54.252 ± 9.843
8 57.625 ± 7.638 60.661 ± 4.991 56.31 ± 09.57 56.888 ± 11.139 61.3 ± 7.961
16 60.93 ± 4.997 63.893 ± 6.62 64.52 ± 08.93 65.907 ± 3.691 69.28 ± 2.358
Airline (3) 4 46.175 ± 4.922 44.167 ± 10.752 42.76 ± 13.50 36.243 ± 22.607 50.987 ± 4.936
8 48.039 ± 5.677 50.148 ± 13.429 38.00 ± 17.06 44.972 ± 22.584 55.209 ± 6.049
16 51.754 ± 2.954 54.8 ± 10.49 58.01 ± 08.23 29.238 ± 17.494 60.247 ± 4.577
Emotion (13) 4 8.474 ± 1.028 8.953 ± 2.052 09.20 ± 3.22 8.662 ± 6.213 9.706 ± 1.108
8 8.418 ± 0.792 10.857 ± 3.436 08.21 ± 2.12 8.16 ± 3.266 8.041 ± 2.797
16 8.758 ± 0.854 11.479 ± 2.96 13.43 ± 2.51 8.115 ± 3.66 10.919 ± 1.615
Rating Books (3) 4 39.324 ± 4.985 37.715 ± 5.801 39.42 ± 07.22 25.562 ± 15.207 42.116 ± 4.725
8 38.447 ± 4.07 38.518 ± 5.327 39.55 ± 10.01 34.026 ± 14.123 42.156 ± 4.608
16 41.107 ± 2.683 44.694 ± 7.797 43.08 ± 11.78 32.509 ± 16.132 46.513 ± 3.036
Political Message (9) 4 15.044 ± 1.556 13.888 ± 2.076 15.64 ± 2.73 12.438 ± 1.799 14.421 ± 1.095
8 15.666 ± 1.053 16.155 ± 2.316 13.38 ± 1.74 15.08 ± 2.925 16.919 ± 1.756
16 14.844 ± 1.426 18.324 ± 2.011 20.67 ± 3.89 13.121 ± 3.294 18.319 ± 1.74
Scitail (2) 4 53.66 ± 4.594 50.227 ± 5.69 58.53 ± 09.74 52.296 ± 4.366 54.101 ± 3.759
8 53.212 ± 3.028 54.196 ± 6.678 57.93 ± 10.70 55.964 ± 5.705 56.341 ± 5.786
16 54.53 ± 4.399 57.744 ± 5.696 65.66 ± 06.82 59.675 ± 4.033 59.692 ± 4.227
Sentiment DVD (2) 4 53.78 ± 1.316 51.06 ± 3.302 54.98 ± 3.96 56.06 ± 2.408 55.003 ± 2.936
8 53.8 ± 2.843 55.19 ± 3.298 55.63 ± 4.34 56.98 ± 3.299 57.527 ± 3.562
16 54.05 ± 1.659 59.45 ± 3.84 58.69 ± 6.08 58.95 ± 2.813 60.76 ± 2.944
Sentiment Books (2) 4 57.56 ± 4.571 55.53 ± 4.097 54.81 ± 3.75 59.89 ± 5.385 58.67 ± 4.753
8 60.76 ± 4.177 58.97 ± 4.909 53.54 ± 5.17 64.34 ± 2.565 64.78 ± 2.615
16 60.98 ± 2.857 65.5 ± 7.026 65.56 ± 4.12 66.36 ± 2.183 67.453 ± 3.085
Restaurant (8) 4 48.194 ± 4.881 18.161 ± 2.822 49.37 ± 4.28 24.932 ± 17.102 47.634 ± 5.237
8 57.36 ± 3.681 32.146 ± 5.785 49.38 ± 7.76 29.787 ± 9.573 55.912 ± 4.494
16 64.706 ± 2.913 40.435 ± 3.348 69.24 ± 3.68 29.154 ± 13.537 61.716 ± 2.208
CoNLL (4) 4 45.104 ± 7.724 35.438 ± 7.324 50.44 ± 08.57 27.02 ± 7.346 52.724 ± 5.84
8 46.897 ± 4.183 44.259 ± 4.886 50.06 ± 11.30 31.296 ± 17.487 60.374 ± 3.731
16 53.969 ± 3.252 52.116 ± 5.354 74.47 ± 03.10 22.923 ± 7.933 67.496 ± 4.551
Table 1: Classification performance (accuracy) of our methods (SLP and DeepSLP) and baselines. Entries in green indicate the best model; * is the baseline as reported in Bansal et al. (2019). Subscripts for SLP and DeepSLP refer to the (non-fine-tuned) encoder used.
Our results (Table 1) demonstrate that DeepSLPBERT outperforms BERTfine-tuned in 31/48 tasks, SLPBERT in 43/48 tasks and 1-NN in 41/48 tasks, demonstrating the usefulness of soft-label prototypes and superiority over the “standard” LLM fine-tuning paradigm, as well as the simpler SLP variant. SLPBERT, on the other hand, fares worse than BERTfine-tuned and 1-NN, outperforming the former in only 19/48 tasks and the latter in 17/48 tasks, while exhibiting high standard deviations. Overall, DeepSLP is the best performing method, demonstrating the highest accuracy in 31/48 tasks, while being on-par with the second best model (BERTfine-tuned) on the remaining tasks (15/48 tasks).
SLP vs. DeepSLP
From Table 1, we can see that SLP exhibits high standard deviations, in particular for tasks with a high number of classes ($N>2$). [^5] The only exceptions to this are Disaster ($N=2$), where we observe a high standard deviation across all models; and Emotion ($N=13$) and Political Message ($N=9$), where we observe lower standard deviations. There we only find lines containing 2 centroids (except for one with 3 centroids). This low number of classes per line is therefore responsible for the higher stability and, in turn, lower standard deviations. SLP’s behaviour can be explained by the fact that it can, at times, fail to select the nearest centroid (proof provided in Theorem 12.1), as demonstrated in the example in Figure 5. If we consider the case where $N=2$, then SLP essentially acts as a 1-NN with soft labels trivially at $[1,0]$ and $[0,1]$, with class centroids acting as the nearest neighbour. However, when generalising beyond this setting, the model’s stability is affected. SLP optimises soft labels using the geometric properties of a line and does not consider each (training) data point individually – the soft labels produced by SLP are constants. DeepSLP, on the other hand, learns from training data and produces soft labels as a function of an input point; therefore, it has the ability to output soft labels based on the location of an input (test) point (with the location of the prototypes being fixed).
Figure 5: The centroids for classes A (green), B (pink) and C (red) are located at $(-12,0)$, $(5,0)$ and $(15,0)$ respectively. Unlike most Euclidean distance based algorithms such as 1-NN, the test point marked by $x$ is assigned to B (instead of the nearest class, A) with SLP. The soft labels at $(-15,0)$ and $(15,0)$ are represented in the pie charts on the right.
# SLP DeepSLP
1
$015.6422e-$ 01
$019.7887e-$ 01
$014.3577e-$ 01
$041.3500e-$ 04
$169.7973e-$ 16
$022.0995e-$ 02
2
$135.3090e-$ 13
$022.5240e-$ 02
$014.3212e-$ 01
$051.0000e-$ 05
$015.6787e-$ 01
$019.7475e-$ 01
Table 2: Soft labels derived from SLP and DeepSLP. # denotes the index of the soft-label prototype lying on the line.
In Table 2, we further present the probability distribution of two soft-label prototypes for the category Airline using SLP and DeepSLP with the 16-shot setting. It can be seen that the distribution of soft labels derived from SLP tends to be more skewed, with the probability distribution lying heavily on 2/3 classes, which is in contrast to DeepSLP that seems to be more uniform. This is a pattern that we generally observe across all tasks. This can further explain the lower performance of SLP, as this behaviour essentially tends to dismiss an entire class. DeepSLP, on the other hand, leads to better generalisation and a more stable performance in our highly limited data setting.
Number of shots
For DeepSLP, we observe a general trend where performance increases as the number of shots increases from 4 to 16. On the other hand, for SLP, we do not see as clear a picture, indicating that more data does not necessarily improve performance (e.g., there is a decrease in performance for SLP in $7/16$ test tasks as we move from 8 to 16 shots).
(Deep)SLP training time and model parameters
Training time reduces to three operations: encoding the data, generating lines from class centroids, and constructing soft-label prototypes at the ends of the lines. Generally, we observe that (a) all lines are constructed under a second, (b) soft labels via optimisation of linear constraints using the MOSEK solver are generated in less than a second, and (c) soft labels learned by minimising training loss require less than 10 seconds per line (tasks with $N=2$ or $N=3$ use only one line, except for emotion which has $N=13$ and takes less than a minute). Furthermore, the number of parameters are substantially smaller compared to those in fine-tuning an LLM – we only need to fine-tune $1500-10000$ parameters (based on the number of classes) for the linear layer of each soft-label prototype for DeepSLP, compared to $>10^{8}$ parameters for finetuning BERT. DeepSLP does not require a GPU.
9 Further analysis
9.1 DeepSLP architecture analysis
[…]
9.2 Using a more powerful encoder
While our focus is on extreme few-shot learning utilising limited amounts of training data, in this section, we seek to analyse model behaviour when auxiliary labelled data is available. Specifically, we utilise another encoder, MT-DNN Liu et al. (2019), a BERT model that has been fine-tuned on the GLUE Wang et al. (2018) dataset (similarly to Bansal et al. (2019)), and which we further fine-tune on our data, similarly to our fine-tuned BERT baseline. In Table 4, we present the performance of MT-DNN (fine-tuned) and a DeepSLP variant that uses the MT-DNN encoder (fixed). We can see that the performance of the two is closer in this case (compared to when using the BERT encoder in Table 1) – DeepSLP outperforms MT-DNN in
Category (Classes) Shot BERT DeepSLPBERT
Rating DVD (3) 4 36.142 ± 8.17 39.566 ± 5.086
8 37.855 ± 5.208 38.788 ± 4.449
16 41.465 ± 4.597 40.53 ± 4.375
Rating Electronics (3) 4 37.815 ± 9.3 39.977 ± 5.959
8 40.901 ± 3.97 41.926 ± 3.985
16 45.289 ± 2.832 44.917 ± 3.164
Political Bias (2) 4 52.466 ± 2.776 53.251 ± 4.042
8 57.819 ± 4.709 58.209 ± 5.198
16 61.025 ± 2.879 61.479 ± 2.974
Rating Kitchen (3) 4 38.954 ± 12.067 39.624 ± 6.787
8 40.511 ± 8.019 41.081 ± 6.777
16 44.923 ± 3.536 45.801 ± 4.562
Political Audience (2) 4 51.069 ± 1.259 51.741 ± 2.827
8 54.117 ± 2.824 54.506 ± 3.274
16 56.827 ± 3.438 56.956 ± 3.045
Sentiment Kitchen (2) 4 58.697 ± 4.943 60.76 ± 4.426
8 63.473 ± 3.207 65.733 ± 3.198
16 69.14 ± 2.413 69.18 ± 2.589
Disaster (2) 4 54.005 ± 10.75 54.252 ± 9.843
8 61.644 ± 6.371 61.3 ± 7.961
16 68.378 ± 2.686 69.28 ± 2.358
Airline (3) 4 44.714 ± 14.082 50.987 ± 4.936
8 53.887 ± 7.074 55.209 ± 6.049
16 59.967 ± 3.842 60.247 ± 4.577
Emotion (13) 4 10.068 ± 2.714 9.706 ± 1.108
8 11.876 ± 0.743 8.041 ± 2.797
16 12.124 ± 1.945 10.919 ± 1.615
Rating Books (3) 4 40.832 ± 8.488 42.116 ± 4.725
8 41.842 ± 5.425 42.156 ± 4.608
16 46.494 ± 2.465 46.513 ± 3.036
Political Message (9) 4 15.311 ± 2.265 14.421 ± 1.095
8 18.244 ± 1.636 16.919 ± 1.756
16 18.971 ± 2.206 18.319 ± 1.74
Scitail (2) 4 51.474 ± 6.275 54.101 ± 3.759
8 55.173 ± 6.254 56.341 ± 5.786
16 59.887 ± 3.667 59.692 ± 4.227
Sentiment DVD (2) 4 54.457 ± 2.97 55.003 ± 2.936
8 56.54 ± 4.009 57.527 ± 3.562
16 60.593 ± 2.73 60.76 ± 2.944
Sentiment Books (2) 4 57.217 ± 3.822 58.67 ± 4.753
8 63.87 ± 3.982 64.78 ± 2.615
16 67.137 ± 3.107 67.453 ± 3.085
Restaurant (8) 4 25.63 ± 3.834 47.634 ± 5.237
8 37.475 ± 2.683 55.912 ± 4.494
16 44.132 ± 3.33 61.716 ± 2.208
CoNLL (4) 4 47.97 ± 5.926 52.724 ± 5.84
8 62.647 ± 4.582 60.374 ± 3.731
16 70.314 ± 2.23 67.496 ± 4.551
Table 3: Contrasting the performance of DeepSLP to a counterpart (BERT) that only tests the architecture of DeepSLP.
23/48 tasks, but the differences are small ($<1\%$ in 18/48 tasks, with similar standard deviation). Again, this highlights DeepSLP’s effectiveness in data lean settings. Furthermore, DeepSLP does not require a GPU while only trains a fraction of the parameters used in MT-DNN finetuning.
Category (Classes) Shot MT-DNN DeepSLPMT-DNN
Rating DVD (3) 4 51.26 ± 8.3 57.87 ± 7.3
8 55.39 ± 10.07 53.96 ± 5.68
16 60.02 ± 1.8 56.2 ± 6.43
Rating Electronics (3) 4 59.82 ± 4.24 57.63 ± 7.91
8 63.37 ± 0.78 60.51 ± 2.91
16 62.85 ± 0.62 61.47 ± 2
Political Bias (2) 4 62.46 ± 4.09 61.25 ± 9.25
8 66.37 ± 0.36 63.83 ± 4.53
16 66.10 ± 3.79 65.63 ± 2.03
Rating Kitchen (3) 4 54.80 ± 13.02 56.81 ± 13.42
8 57.08 ± 11.95 57.47 ± 9.99
16 63.95 ± 4.34 62.02 ± 4.83
Political Audience (2) 4 51.03 ± 7.22 52.87 ± 7.28
8 56.74 ± 5.30 56.78 ± 3.72
16 58.70 ± 2.15 58.78 ± 2.83
Sentiment Kitchen (2) 4 82.50 ± 3.49 86.65 ± 0.54
8 86.67 ± 0.60 86.09 ± 1.26
16 86.63 ± 0.50 86.87 ± 0.26
Disaster (2) 4 50.29 ± 3.60 51.28 ± 3.44
8 51.61 ± 5.07 56.67 ± 5.29
16 56.08 ± 4.77 53.23 ± 1.94
Airline (3) 4 70.41 ± 2.45 63.65 ± 11.7
8 70.22 ± 2.47 70.8 ± 2.69
16 71.04 ± 0.64 72.51 ± 1.14
Emotion (13) 4 12.88 ± 0.73 10.39 ± 2.71
8 15.39 ± 1.37 11.17 ± 2.16
16 17.45 ± 0.56 11.24 ± 1.99
Rating Books (3) 4 65.52 ± 8.96 62.25 ± 6.79
8 69.28 ± 0.27 63.27 ± 3.16
16 68.94 ± 0.78 68.02 ± 1.8
Political Message (9) 4 19.78 ± 1.10 18.38 ± 2.13
8 21.56 ± 0.65 21.16 ± 1.72
16 23.63 ± 2.47 23.02 ± 1.11
Scitail (2) 4 56.11 ± 3.52 63.71 ± 6.85
8 69.36 ± 3.41 73.33 ± 5.63
16 74.18 ± 3.38 73.81 ± 3.66
Sentiment DVD (2) 4 84.30 ± 0.96 84.57 ± 1.05
8 84.90 ± 0.75 84.18 ± 1.41
16 84.73 ± 0.52 85.09 ± 0.28
Sentiment Books (2) 4 85.53 ± 0.99 86.23 ± 0.4
8 85.43 ± 0.34 85.6 ± 1.23
16 85.73 ± 0.71 86.18 ± 0.33
Restaurant (8) 4 28.37 ± 2.65 34.28 ± 5.23
8 46.65 ± 3.80 44.15 ± 5.43
16 58.66 ± 1.24 47.09 ± 3.71
CoNLL (4) 4 35.97 ± 2.38 47.33 ± 6.2
8 49.04 ± 0.63 59.13 ± 5.14
16 64.55 ± 1.14 61.96 ± 5
Table 4: Contrasting the performance of DeepSLP to MT-DNN.
10 Ensemble properties of DeepSLP
In this study, we compare and contrast DeepSLP to ensembles and draw similarities between the two, shedding further light into the effectiveness of our approach. Each prediction decision by DeepSLP is the result of two soft-label prototypes – those that lie on each end of the line nearest to a test point $x$. An analogy can then be drawn between the prototypes used at prediction time and those individual (albeit independent) models that are utilised by an ensemble when producing the final classification. While DeepSLP prototypes are not independent but are rather trained jointly (and share the same encoder), in what follows, we demonstrate that they display several properties of ensemble methods, while being computationally efficient and utilising a small number of parameters. For the analyses below, we consider the tasks Airline and Disaster using an 8-shot setting and evaluate on the test data for each.
10.1 Individual vs joint prediction
In a similar way as an ensemble exhibits superior performance to the individual models it utilises, we seek to assess whether the joint utilisation of prototypes at prediction time is indeed more effective than utilising each prototype individually. To evaluate this, we measure the probability distribution of each setting on the test data using negative log-likelihood:
$NLL(\mathbf{f(x)},y)\triangleq-log(\mathbf{f^{(y)}(x)})$
Following Abe et al. (2022), for a strictly convex function such as $NLL$, we use Jensen’s inequality:
$NLL(\mathbf{F(x)},y)\;\leq\;\mathbb{E}[NLL(\mathbf{f(x)},y)]$
where $\mathbf{F(x)}$ is the ensemble and $\mathbf{f(x)}$ are the constituent models. The idea is that the probability distribution of the ensemble fits the target distribution more closely than the corresponding expected probability distributions of its constituent models. For joint soft-label prototypes parameterised by $g_{1}$ and $g_{2}$ and located at $p_{1}$ and $p_{2}$, we have:
$NLL\left(softmax\left(\frac{g_{1}(f(x))}{||f(x)-p_{1}||})+\frac{g_{2}(f(x))}{||f(x)-p_{2}||}\right),y\right)\;\;\;\;\;$
$=-\sum log\left(softmax\left(\frac{g_{1}^{y}(f(x))}{||f(x)-p_{1}||})+\frac{g_{2}^{y}(f(x))}{||f(x)-p_{2}||}\right)\right)\;$
For the individual soft-label prototypes, weighing the outputs by distance does not change the final softmax probability distribution; therefore, we can define $\mathbb{E}[NLL(\mathbf{f(x)},y)]$ as their average:
$-\frac{1}{2}\sum log(softmax(g_{1}^{y}(f(x))+log(softmax(g_{2}^{y}(f(x))$
We plot the negative log likelihoods for Airline and Disaster on the test set after finetuning each model on ten subsets of few-shot training data (as explained before in Section 5) in Figure 6 to assess whether DeepSLP exhibits this property of ensemble methods.
Figure 6: $NLL(\mathbf{F(x)},y)$ vs. $\mathbb{E}[NLL(\mathbf{f(x)},y)]$ for Disaster (top) and Airline (bottom). The ith subscript refers to the ith soft-label prototype.
We find that $NLL(\mathbf{F(x)},y)$ is much lower than $\mathbb{E}[NLL(\mathbf{f(x)},y)]$, which confirms that the joint utilisation of prototypes results in better predictions than if they were to be used individually. Our experiments confirm that this is a general trend we observe across tasks. In the following sections, we investigate the reasons behind the high values observed for $\mathbb{E}[NLL(\mathbf{f(x)},y)]$, and compare the jointly trained prototypes against a strong baseline, the fine-tuned BERT baseline.
10.2 Jointly utilised soft-label prototypes improve diversity
Diversity in ensemble classifications refers to the difference in the probability distribution on out-of-distribution (ood) data for classifications between individual models and the ensemble. We use this definition to ascertain the diversity of classifications provided by the jointly utilised soft-label prototypes. Diversity is a desirable property as ensemble predictions are generally more robust due to diversity between the predictions of their individual members Lee et al. (2015).
Existing work Ashukha et al. (2020); Lakshminarayanan et al. (2017) defines ensemble uncertainty as the sum of ensemble diversity and the expected average model uncertainty on ood data. Based on Abe et al. (2022), it is calculated as:
$H([y|F(x)]=\frac{-1}{C}\sum p(y_{i}|F(x))log(p(y_{i}|F(x)))$
If we use the Jenson-Shannon divergence as a diversity measure for an ensemble given by $JSD_{p(f)}[y|f(x)]=\frac{1}{M}\sum KL[y|f(x)||y|F(x)]$ where KL is the average KL divergence between the output distribution of each soft-label prototype and the jointly utilised soft-label prototypes, from Abe et al. (2022), this expression reduces to:
$H([y|F(x)]=\stackrel{{\scriptstyle\underline{ens.\;diversity}}}{{JSD_{p(f)}[y|f(x)]}}+\stackrel{{\scriptstyle\underline{avg.\;model\;uncert.}}}{{E_{p(f)}[H[y|f(x)]]}}$
We contrast ensemble uncertainty and single model uncertainty using the fine-tuned BERT model for the task airline in Figure 7, but note that similar trends are observed across all tasks. We note that the uncertainty of jointly utilised soft-label prototypes is generally higher than that of the fine-tuned BERT model. As the average model uncertainty of individual soft-label prototypes is negligibly low, the uncertainty in the joint case is driven mainly by the diversity of the ensemble. This is in line with previous work which attributes an increase in uncertainty in ensembles due to diversity (Lakshminarayanan et al., 2017; Dietterich, 2000; Wilson and Izmailov, 2020). Though not strictly an ensemble, our approach exhibits similar properties (higher uncertainty driven by model diversity), as well as a general reduction in standard deviation compared to fine-tuned BERT.
Figure 7: Ensemble uncertainty contrasted against the uncertainty of fine-tuned BERT, where we observe that DeepSLP’s uncertainty $F(x)$ (given by yellow) is driven by ensemble diversity, given by $JSD(F(x))$ in blue.
The above provide evidence that our approach as a whole exhibits desirable properties of ensembles which drive a higher performance but which do not lead to a higher training time nor compute.
11 Conclusion
We developed a novel few-shot learning approach for data lean settings that is based on soft-label prototypes capturing the simultaneous membership of data points over several classes. We presented two variants, one that generates fixed prototypes via solving for linear constraints, and another that learns and generates them dynamically within a neural gradient-based optimisation approach. We evaluated our methods on 48 different tasks / settings and experimentally demonstrated that our neural approach outperforms a range of strong baselines. Finally, we presented an in-depth analysis of the characteristics of our approach and analysed the reasons behind its effectiveness.